# Inverse 3D microscopy rendering for cell shape inference with active mesh

### v2

@@ 5.3 PSF optimization step
PSF parameter optimization aims to accurately reconstruct both bright and dark regions of the image. The L2 norm minimization, $\phi_{0}=||u-m||^{2}$ is well-suited for this task. We optimize the covariance matrix $\Sigma$ using Adam [35].
@@ 6.1 Benchmarks on artificial images
We first benchmark our method on artificial images generated from common computer graphics meshes and a simulated dividing cell (figure 6. Images are rendered using a Gaussian isotropic PSF, which was also used for inference to isolate the effect of shape reconstruction. Starting from elementary mesh shapes (a sphere or torus, depending on topology), we perform 10,000 optimization iterations per example. Despite the absence of remeshing or collision detection algorithms [9, 45], our regularized shape optimization accurately converge to the original shapes.
@@ 6.2 Robustness to initial conditions and noise
Then we evaluate the convergence robustness to initial conditions and noise. As a metric, we compute the mean-squared error (MSE) between the reconstructed shape—rendered as an artificial image—and the target image generated from the initial mesh. Using Spot (cow) as the reference mesh, we assess the MSE when the initial mesh was scaled (zoomed in/out) and shifted from its barycenter (Figure 7, top and middle). While our method is generally robust to initialization, large shifts from the target increase the number of iterations needed for convergence. Notably, our approach demonstrates high robustness to random noise added to the image (Figure 7, bottom), which is particularly advantageous for microscopy applications.
Figure 8: Cell shape inference in C. elegans embryos. Starting from a foam ground truth, we optimize both vertices position and PSF parameters to recover individual cell shapes from real confocal microscopy data (see also supplementary videos 5-7).
6.3 Comparison with active mesh algorithm
We compare our method to DM3D (Deforming Mesh 3D), a recent active mesh algorithm designed for detecting cell interfaces in fluorescence microscopy images [68]. Traditional active mesh methods were primarily developed for bright bulk regions [11, 16] rather than structures defined by a bright intensity band. Using the dataset provided by the authors, which includes 15 mouse organoid 3D images and their corresponding inferred cellular meshes, we apply our method to their original images. As evaluation metrics between DM3D and deltaMic, we compute the MSE, Structural Similarity (SSIM) [75], and Peak-Signal-to-Noise Ratio (PSNR) [27] between the reconstructed shapes from both methods, rendered as artificial images using our inferred PSF.
@@ 6.4 Shape inference from embryo microscopy
Finally, we applied our method to infer cell shapes in early-stage C. elegans embryos, using 3D confocal microscopy images from [10] of size $205\times 285\times 134$. Cell clusters were represented as non-manifold multimaterial meshes [8, 49, 19], and we employed a multimaterial mesh-based surface tracking method [13] to handle remeshing, collision detection, and topology transitions. As an initial guess, we generated a foam-like multimaterial mesh with the correct number of cells. The resulting cell clusters, containing $\approx 4000-10000$ vertices, were optimized to fit the microscopy data. We jointly optimized the PSF and the mesh, following the strategy in Section 5.1. After 10,000 optimization steps, our method accurately reconstructed cell shapes in 4-, 6-, and 16-cell embryos, producing rendered images that closely matched the original microscopy images (Figure 8).
@@ 7 Conclusion
We introduced deltaMIC, a differentiable 3D microscopy image renderer that generates realistic 3D confocal fluorescence images from a surface mesh and a parametrized PSF. Our GPU-parallelized mesh FT implementation enables efficient forward and backward computations, further optimized with a spectral narrow-band method. deltaMIC is relatively robust to random noise in the image and mesh initialization, accurately recovering fine details of complex shapes from artificial and real 3D microscopy images thanks to its ability to fine-tune the shape loss function via the PSF. Unlike traditional active mesh methods, it does not require fine-tuned shape regularization terms to achieve smooth results. However, knowledge-based priors can be incorporated to constrain cell shapes, opening the avenue for direct implementation of inverse mechanical problems from microscopy data [28] through integration with a differentiable physical simulator. Our approach promises numerous other practical and theoretical extensions. As an active contour method, it has the potential to natively allow for biological shape tracking in timelapse microscopy. Implementing more photorealistic PSF models based on specific microscope modalities [23, 24, 40] could allow performing blind deconvolution of microscopy images with shape priors [42]. Regarding the shapes, natural extensions of the method to 1D objects, such as cytoskeletal filaments whose segmentation remain a major challenge [60] or 3D objects such nuclei or condensates, could pave the way for active multidimensional mesh algorithms. From an implementation perspective, further acceleration strategies could be explored to decrease even further the computational cost bottleneck for large 3D images. Finally, through its inherent differentiability, the method can be easily integrated as building block in a larger (deep)-learning pipeline [34, 53].
@@ 8 Acknowledgments
SI, FB and HT are supported by the CNRS and Collège de France. HT received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 949267). SI was funded by Ecole Polytechnique (AMX grant). We thank Baptiste Nicolet for insightful disussions, and H. Borja da Rocha and K. Crane for their meshes.

### v3

@@ 5.2 PSF optimization step
PSF parameter optimization aims to accurately reconstruct both bright and dark regions of the image. A regular $L^{2}$ norm minimization, $\mathcal{L}_{0}=||I-I_{\text{GT}}||^{2}$ is well-suited for this task. In this step, we optimize the covariance matrix $\Sigma$ using the Adam optimizer [43].
@@ 6.1 Benchmarks on artificial images
We first benchmark our method on artificial images generated from common computer graphics meshes and a simulated dividing cell (Figure 7). Images are rendered using a Gaussian isotropic PSF, which was also used for inference to isolate the effect of shape reconstruction. Starting from elementary mesh shapes (a sphere or torus, depending on topology), we perform 10,000 optimization iterations per example. Despite the absence of remeshing or collision detection algorithms [10, 52], our regularized shape optimization accurately converges to the original shapes.
Figure 7: Shape inference from 3D artificial images: Starting from spherical or toroidal meshes, the optimization process reconstructs meshes that closely align with the original shapes (see also: Supplemental videos 1–4).
@@ 6.2 Robustness to initial conditions and noise
Next, we evaluate the convergence robustness to initial conditions and noise. As a metric, we compute the mean-squared error (MSE) between the reconstructed shape—rendered as an artificial image—and the target image generated from the initial mesh. Using Spot (cow) as the reference shape, we assess the MSE when the starting mesh is scaled (zoomed in/out) and shifted from its barycenter (Figure 8, top and middle). While our method is generally robust to initialization, large shifts from the target increase the number of iterations needed for convergence. Notably, our approach demonstrates high robustness to random noise added to the image (Figure 8, bottom), which is particularly advantageous for microscopy applications.
Figure 9: Benchmarking deltaMic vs DM3D [75] on 3D mouse organoid images.
Figure 10: Cell shape inference in early embryos: Starting from a foam-like mesh, both vertex positions and PSF are optimized to infer individual cell shapes in 4-cell mouse (top), 8-cell ascidian (middle), and 16-cell worm (bottom) embryos, using 3D fluorescence images acquired respectively with single-view, dual-view light-sheet, and point-scanning microscopes (see also: Supplemental videos 5–9). Mouse and ascidian data - courtesy of K. Cavanaugh and D. Gonzalez-Suarez.
6.3 Benchmarking against an active mesh method
Traditional active mesh methods were primarily developed for segmenting bright volumetric regions [12, 21], making them less suited for structures characterized by narrow intensity bands. We benchmark our approach against DM3D (Deforming Mesh 3D) [75], a recent active mesh algorithm designed to detect cell interfaces in fluorescence microscopy images. Using the authors’ dataset of 15 mouse organoid 3D images with corresponding inferred meshes, we reapply deltaMic to the original image volumes. We then compare the outputs of DM3D and deltaMic using mean squared error (MSE), structural similarity index (SSIM) [82], and peak signal-to-noise ratio (PSNR) [33], computed between the reconstructed shapes rendered as synthetic images with our inferred PSF. Our method consistently outperforms DM3D across all metrics (Figure 9).
@@ 6.4 Shape inference from embryo microscopy
We applied our method to infer cell shapes in early-stage embryos from different species and 3D fluorescence microscopy modalities: mouse (single-view light-sheet), ascidian (dual-view light-sheet), and worm (point-scanning confocal) [11]. Cell clusters were represented as non-manifold multimaterial meshes [9, 56, 25], and we used a multimaterial mesh-based surface tracking method [15] to handle remeshing, collision detection, and topological transitions. As an initial guess, we generated a foam-like multimaterial mesh with the correct number of cells. The resulting cell clusters, ranging between 1000–10000 vertices, were optimized to fit the microscopy data. We jointly optimized the PSF and mesh following the strategy in Section 5.1. deltaMic accurately reconstructed cell shapes in 4-, 8-, and 16-cell embryos, converging within 1k and 10k optimization steps, yielding rendered images that closely matched the original microscopy data (Figure 10).
@@ 7 Conclusion
We introduced deltaMic, a first-of-its-kind differentiable 3D microscopy image creator that renders realistic 3D confocal fluorescence images from a surface mesh and a parametrized PSF. Our GPU-parallelized mesh FT implementation enables efficient forward and backward computations, further optimized with a spectral narrow-band method. deltaMic is relatively robust to random noise in the image and mesh initialization, accurately recovering fine details of complex shapes from artificial and real 3D microscopy images thanks to its ability to fine-tune the shape loss function via the PSF. Unlike traditional active mesh methods, it does not require fine-tuned shape regularization terms to achieve smooth results. However, knowledge-based priors can be incorporated to constrain cell shapes, opening the avenue for direct implementation of inverse mechanical problems from microscopy data [34] through integration with a differentiable physical simulator.
Our approach opens numerous avenues for both practical applications and theoretical extensions. As an active contour method, it naturally lends itself to biological shape tracking in time-lapse microscopy. Incorporating more photorealistic PSF models tailored to specific microscopy modalities [29, 30, 47, 1] could enable blind deconvolution of microscopy images with integrated shape priors [49, 17]. But extending our framework to handle space-variant PSFs [22, 16], which violate translational invariance, may require significant theoretical and computational developments. In terms of geometry, the method could be adapted to handle 1D structures—such as cytoskeletal filaments, whose segmentation remains a major challe [67]—or extended to fully 3D shapes like nuclei or condensates, paving the way for multidimensional active mesh algorithms. On the implementation side, additional acceleration strategies could help mitigate the computational demands of large 3D volumes. Finally, thanks to its full differentiability, our method can be readily integrated as a module within larger (deep-)learning pipelines [39, 60].
@@ 8 Acknowledgments
SI, AS, FB and HT are supported by the CNRS and Collège de France. HT received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 949267) and from the French Agence Nationale de la Recherche (Grants ANR-22-CE13-0036 and ANR-23-CE30-0013). SI fellowship was funded by Ecole Polytechnique (AMX grant). We thank Baptiste Nicolet for insightful discussions, H. Borja da Rocha and K. Crane for sharing meshes, K. Cavanaugh and D. Gonzalez Suarez for sharing fluorescence microscopy images of mouse and ascidian embryos.